# NeuralFuse: Learning to Recover the Accuracy of Access-Limited Neural Network Inference in Low-Voltage Regimes

### v2

@@ 4.5 Model Size and Efficiency of NeuralFuse
To provide a full performance characterization of NeuralFuse, we analyze the relationship between the final recovery of each base model and generators of varying parameter counts. The efficiency ratio is defined as the recover percentage in perturbed accuracy divided by the parameter count of NeuralFuse. We compare the efficiency ratio for all NeuralFuse generators (trained on CIFAR-10) in Table 3. The results show that UNet-based generators have better efficiency than both Convolution-based and Deconvolution-based ones per million parameters.
![[Uncaptioned image]](x19.png)
![[Uncaptioned image]](x20.png)
Figure 4: Reduced-precision accuracy
4.6 NeuralFuse on Reduced-precision Quantization
@@
Here we explore the robustness of NeuralFuse to low-precision quantization on model weights. Uniform quantization is the de-facto method for quantizing model weights (Gholami et al., 2022). However, it is possible to cause an accuracy drop due to precision loss. As a bit-error-oriented protector, we seek to understand whether NeuralFuse could also make a recovery to mitigate the model’s accuracy drop in this scope. We conducted an experiment that uniformly quantized the model weights to a lower bit precision and reported the resulting accuracy. Specifically, we apply symmetric uniform quantization on our base models with different numbers of bits to induce precision loss. Therefore, the quantized weight (integer) is defined as , where denotes the original model weight (full precision), denotes the quantization scale parameter, and is the precision (number of bits) that used to quantize the models. In addition, we considered the possibility of random bit errors (the low-voltage setup) in low-precision regimes.
@@
We use GTSRB pre-trained ResNet18 as an example to evaluate two NeuralFuse generators (ConvL and UNetL, trained with B.E.R.), and vary the precision from 8 bits to 2 bits (integer). The result is shown in Figure 4. We find that when bits, NeuralFuse can take effect to recover the accuracy in both scenarios. When , while NeuralFuse can still handle the bit-error-free model (top panel), it exhibits a limited ability to recover the random bit-error case (bottom panel). We find the result encouraging because the observed robustness to reduced-precision inference is an emergent ability of NeuralFuse. That is, NeuralFuse was only trained with random bit errors, but it demonstrated high accuracy to unseen bit quantization errors. This experiment showcases NeuralFuse’s potential in protecting against accuracy drops caused by different sources of bit errors. More experimental results of different base models and datasets can be found in Appendix H.
@@ 4.7 Extended Analysis
We highlight some key findings from the additional results in Appendix. In Appendix E, we compare NeuralFuse to the simple baseline of learning a universal input perturbation. We find that the baseline is much worse than NeuralFuse, which validates the necessity of adopting input-aware transformation for learning error-resistant data representations in low-voltage scenarios. In Appendix G, we find that ensemble training of white-box surrogate base models can further improve the transferability of NeuralFuse in the restricted-access setting. In Appendix K and Appendix L, we present visualization results of data embeddings and transformed inputs via NeuralFuse. In Appendix J, we show that NeuralFuse can further recover the accuracy of a base model trained with adversarial weight perturbation in the low-voltage setting.
@@ 5 Conclusion
In this paper, we propose NeuralFuse, the first non-intrusive post-hoc protection module for model inference against bit errors induced by low voltage. NeuralFuse is particularly suited for practical machine deployment settings where access to the base model is limited or relaxed. The design of NeuralFuse includes a novel loss function and a new optimizer named EOPM to handle simulated randomness in perturbed models. Our comprehensive experimental results and analysis show that NeuralFuse can significantly recover test accuracy (by up to ) while simultaneously enjoying up to reduction in memory access energy. Furthermore, NeuralFuse demonstrates high transferability (to access-constrained models), and versatility (e.g., robustness to low-precision quantization). Our results show that NeuralFuse provides significant improvements in mitigating the energy-accuracy tradeoff of neural network inference in low-voltage regimes and sheds new insights on green AI technology. Our future work includes extending our study to other neural network architectures and modalities, such as transformer-based language models.
Impact Statements
We see no immediate ethical or negative societal consequences of our work, and it holds the potential for positive social impacts. Previous works have highlighted scenarios affected by energy concerns, necessitating strategies to reduce energy consumption (Yang et al., 2017, 2019a). By enhancing the accuracy of models in low-voltage regimes, NeuralFuse can function as an effective protector and an efficient energy saver. This technology should be particularly valuable for extending the battery life of devices reliant on batteries.
@@ Appendix B Implementation Details of NeuralFuse Generator
We consider two main goals in designing the NeuralFuse Generator: 1) efficiency (so the overall energy overhead is decreased) and 2) robustness (so that it can generate robust patterns on the input image and overcome the random bit flipping in subsequent models). Accordingly, we choose to utilize an encode-decoder architecture in implementing the generator. The design of ConvL is inspired by Nguyen & Tran (2020), in which the authors utilize a similar architecture to design an input-aware trigger generator, and have demonstrated its efficiency and effectiveness. Furthermore, we attempted to enhance it by replacing the Upsampling layer with a Deconvolution layer, leading to the creation of DeConvL. The UNetL-based NeuralFuse draws inspiration from Ronneberger et al. (2015), known for its robust performance in image segmentation, and thus, we incorporated it as one of our architectures. Lastly, ConvS, DeConvS, and UNetS are scaled-down versions of the model designed to reduce computational costs and total parameters. The architectures of Convolutional-based and Deconvolutional-based are shown in Table 4, and the architecture of UNet-based generators is in Table 5. For the abbreviation used in the table, ConvBlock means the Convolution block, Conv means a single Convolution layer, DeConvBlock means the Deconvolution block, DeConv means a single Deconvolution layer, and BN means a Batch Normalization layer. We use , , and Adam optimizer. For CIFAR-10, GTSRB, and CIFAR-100, we set batch size for each base model. For ImageNet-10, we set for ResNet18, ResNet50 and VGG11, and for both VGG16 and VGG19.
@@
Appendix C NeuralFuse’s Energy-Accuracy Tradeoff
@@
In Table 6, we report the total weight memory access (T.W.M.A.) using SCALE-SIM. We then showed the energy-accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 5.
@@
We can then use the MAC values to further approximate the end-to-end energy consumption of the whole model. Assume that all values are stored on SRAM and that a MAC represents single memory access. The corresponding MACs-based energy saving percentage (MAC-ES, %) can be derived from Eq. 7 (c.f. Sec. 4.4), and results can be found in Table 8. We can observe that most combinations can save a large amount of energy, except that VGG11 with two larger NeuralFuse (ConvL and DeConvL) may increase the total energy. These results are consistent with the results reported in Table 2. In addition, we also showed the MACs-based energy-accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 6.
Base Model ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 16.2 28.7 19.0 26.6 23.2 28.7 ResNet50 24.5 29.8 25.7 28.9 27.4 29.8 VGG11 -21.8 23.9 -11.5 16.0 3.6 23.7 VGG16 5.0 27.3 10.0 23.5 17.4 27.2 VGG19 10.4 28.0 14.4 25.0 20.2 28.0
@@ Appendix D Inference Latency of NeuralFuse
In Table 9, we report the latency (batch_size=1, CIFAR-10/ImageNet-10 testing dataset) of utilizing the different NeuralFuse generators with two different base models, ResNet18 and VGG19. We can see that although NeuralFuse brings a certain degree of extra latency, we deemed it an inevitable tradeoff for reducing energy consumption in our setting. Although latency is not the main focus of this paper, we envision future research designing a more lightweight version of the NeuralFuse module or using model compression techniques on NeuralFuse to reduce latency.
@@ Appendix E Ablation Studies
Here, we study the effect of used in EOPM (Eq. 5). In Figure 7, we report the results for ConvL and ConvS on CIFAR-10 pre-trained ResNet18, under a 1% bit error rate (B.E.R.). The results demonstrate that if we apply larger , the performance increases until convergence. Specifically, for ConvL (Figure 6(a)), larger empirically has a smaller standard deviation; this means larger gives better stability but at the cost of time-consuming training. In contrast, for the small generator ConvS (Figure 6(b)), we can find that the standard deviation is still large even trained by larger ; the reason might be that small generators are not as capable of learning representations as larger ones. Therefore, there exists a tradeoff between the stability of the generator performance and the total training time. In our implementation, choosing or is a good balance.
Tradeoff Between Clean Accuracy (C.A.) and Perturbed Accuracy (P.A.).
@@
| Base Model | C.A. | P.A. | ConvL | |||
| C.A. (NF) | P.A. (NF) | R.P. | ||||
@@
[Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.
@@
[Note] B.E.R.: the bit error rate of the base model, C.A. (%): clean accuracy, UIP: universal input transformation parameter, P.A.(%): perturbed accuracy, and R.P.: total recover percentage of P.A. (UIP) v.s. P.A.
@@ Appendix F Additional Experimental Results on Relaxed Access Settings
We conducted more experiments on Relaxed Access settings to show that our NeuralFuse can protect the models under different B.E.R. The results can be found in Sec. F.1 (CIFAR-10), Sec. F.2 (GTSRB), Sec. F.3 (ImageNet-10), and Sec. F.4 (CIFAR-100). For comparison, we also visualize the experimental results in the figures below each table.
@@ F.1 CIFAR-10
Base Model NF C.A. 1% B.E.R. 0.5% B.E.R. P.A. C.A. (NF) P.A. (NF) R.P. P.A. C.A. (NF) P.A. (NF) R.P. ResNet18 ConvL 92.6 89.8 87.8 1.7 48.8 90.4 87.9 2.2 17.8 ConvS 88.2 59.5 11 20.6 91.7 78.4 8.3 8.3 DeConvL 38.9 89.6 88.5 0.8 49.6 70.1 90.2 90.0 0.2 19.9 DeConvS 12.4 82.9 68.8 6.4 29.9 11.6 84.1 79.9 3.6 9.8 UNetL 86.6 84.6 0.8 45.6 89.7 86.3 2.4 16.2 UNetS 84.4 68.8 6.0 29.8 90.9 80.7 5.8 10.7 ResNet50 ConvL 92.6 85.5 53.2 22 27.1 90.3 86.5 3.2 25.5 ConvS 85.2 34.6 14 8.5 90.8 73.3 8.7 12.3 DeConvL 26.1 87.4 63.3 21 37.2 61.0 89.5 87.2 2.5 26.2 DeConvS 9.4 82.4 42.2 17 16.1 10.3 90.3 75.5 8.1 14.5 UNetL 86.2 75.5 12 49.4 89.9 83.9 3.6 22.9 UNetS 77.3 56.2 19 30.1 89.7 76.1 7.2 15.1 VGG11 ConvL 88.4 89.6 87.2 2.9 45.1 89.8 87.0 1.3 23.3 ConvS 84.9 66.3 7.5 24.1 88.2 74.5 5.7 10.9 DeConvL 42.2 89.3 87.2 2.6 45.0 63.6 89.6 86.9 1.1 23.2 DeConvS 11.6 85.6 68.2 7.1 26.0 9.3 88.3 75.7 4.6 12.1 UNetL 87.1 83.6 1.3 41.4 88.0 82.4 1.8 18.8 UNetS 85.5 72.7 4.6 30.5 88.1 75.8 4.3 12.2 VGG16 ConvL 90.3 90.1 86.0 6.2 50.3 90.2 88.5 0.9 21.9 ConvS 87.4 59.6 12 23.9 89.9 77.8 4.8 11.1 DeConvL 35.7 89.7 85.5 6.8 49.8 66.6 89.7 88.2 1.0 21.4 DeConvS 7.9 86.8 66.5 11 30.8 8.1 90.0 78.4 4.7 11.8 UNetL 87.4 83.4 4.4 47.7 89.0 86.2 1.5 19.6 UNetS 87.4 71.2 8.2 35.5 89.0 80.2 3.5 13.7 VGG19 ConvL 90.5 89.8 77.7 19 41.7 90.4 88.1 1.8 23.9 ConvS 87.3 52.7 17 16.7 89.6 74.5 9.0 10.3 DeConvL 36.0 86.3 78.4 18 42.4 64.2 90.4 88.5 1.4 24.3 DeConvS 12.0 86.5 58.2 18 22.2 12.4 89.7 75.2 8.6 11.0 UNetL 86.3 82.1 4.8 46.0 89.1 85.0 2.7 20.8 UNetS 86.3 66.4 13 30.4 89.2 77.1 7.3 12.9 • [Note] C.A. (%): clean accuracy, P.A. (%): perturbed accuracy, NF: NeuralFuse, and R.P.: total recover percentage of P.A. (NF) v.s. P.A.

### v3

@@ 4.5 Model Size and NeuralFuse Efficiency
To arrive at a full performance characterization of NeuralFuse, we analyzed the relationship between the final recovery achieved by each base model in combination with generators of varying parameter counts. For this purpose, we defined efficiency ratio as the recovery percentage in PA divided by NeuralFuse’s parameter count. Table 3 compares the efficiency ratios of all NeuralFuse generators trained on CIFAR-10. Those results show that UNet-based generators had better efficiency per million parameters than either convolution-based or deconvolution-based ones.
4.6 NeuralFuse’s Robustness to Reduced-precision Quantization
@@
Lastly, we also explored NeuralFuse’s robustness to low-precision quantization on model weights. Uniform quantization is the usual method for quantizing model weights (Gholami et al., 2022). However, it is possible for it to cause an accuracy drop due to lack of precision. Given our aim of offering protection from bit errors, we hoped to understand whether NeuralFuse could also recover model-accuracy drops caused by this phenomenon. We therefore uniformly quantized the model weights to a lower bit precision and measured the resulting accuracy. Specifically, we applied symmetric uniform quantization to our base models with various numbers of bits to induce precision loss, and defined the quantized weight (integer) as , where denotes the original model weight (full precision), is the quantization scale parameter, and is the precision (number of bits) used to quantize the models. Bit errors induced by low voltage operation as previously described, are also applied to low precision weights.
@@
We used the GTSRB pre-trained ResNet18 as our example in an evaluation of two NeuralFuse generators, i.e., ConvL and UNetL trained with BER, and varied precision from 8 bits to 2 bits (integer). The results, shown in Figure 4, indicated that when bits, NeuralFuse could effectively recover accuracy in both the low-voltage and low-precision scenarios. When , while NeuralFuse could still handle the bit-error-free model (Fig. 4 top), it exhibited a limited ability to recover the random bit-error case (Fig. 4 bottom). We find these results encouraging, insofar as NeuralFuse was only trained on random bit errors, yet demonstrated high accuracy in dealing with unseen bit-quantization errors. Further experimental results derived from other base models and datasets can be found in Appendix H.
@@ 4.7 Extended Analysis
Here, we would like to highlight some key findings from the additional results in the Appendices. In Appendix E, we compare NeuralFuse against a simple baseline of learning a universal input perturbation. We found that such baseline performed much worse than NeuralFuse at that task, validating the necessity of adopting input-aware transformation if the goal is to learn error-resistant data representations in low-voltage scenarios. In Appendix G, we report that ensemble training of white-box surrogate base models could further improve the transferability of NeuralFuse in restricted-access scenarios. Appendices K and L present visualization results of NeuralFuse’s data embeddings and transformed inputs. In Appendix J, we show that NeuralFuse can further recover the accuracy of a base model trained with adversarial weight perturbation in a low-voltage setting.
@@ 5 Conclusion
In this paper, we have proposed NeuralFuse, the first non-intrusive post hoc module that protects model inference against bit errors induced by low voltage. NeuralFuse is particularly well suited to practical machine-deployment cases in which access to the base model is limited or relaxed. The design of NeuralFuse includes a novel loss function and a new optimizer, EOPM, that enable it to handle simulated randomness in perturbed models. Our comprehensive experimental results and analysis show that NeuralFuse can recover test accuracy by up to while simultaneously enjoying an up to reduction in memory-access energy requirements. Moreover, NeuralFuse demonstrates high transferability to access-constrained models and high versatility, e.g., robustness to low-precision quantization. In short, NeuralFuse is a notable advancement in mitigating neural-network inference’s energy/accuracy tradeoff in low-voltage regimes, and points the way to greener future AI technology. Our future work will include extending this study to other neural-network architectures and modalities, such as transformer-based language models.
Limitations. We acknowledge the challenge of achieving significant power savings without accuracy loss and view NeuralFuse as a foundational, proof-of-concept step toward this goal. Future research could enhance this approach by optimizing the pre-processing module to adapt to specific error characteristics of low-voltage SRAM or by integrating lightweight hardware modifications to further improve the energy-accuracy trade-off.
Broader Impacts. We see no ethical or immediate negative societal consequence of our work, and it holds the potential for positive social impacts, from environmental benefits to improved access to technology and enhanced safety in critical applications.
Acknowledgments and Disclosure of Funding
We thank the anonymous reviewers for their insightful comments and valuable suggestions. The research described in this paper was conducted in the JC STEM Lab of Intelligent Design Automation, which is funded by The Hong Kong Jockey Club Charities Trust in support of Tsung-Yi Ho.
@@ Appendix B Implementation Details of NeuralFuse Generator
We consider two main goals in designing the NeuralFuse Generator: 1) efficiency (so the overall energy overhead is decreased) and 2) robustness (so that it can generate robust patterns on the input image and overcome the random bit flipping in subsequent models). Accordingly, we choose to utilize an encode-decoder architecture in implementing the generator. The design of ConvL is inspired by Nguyen and Tran (2020), in which the authors utilize a similar architecture to design an input-aware trigger generator, and have demonstrated its efficiency and effectiveness. Furthermore, we attempted to enhance it by replacing the Upsampling layer with a Deconvolution layer, leading to the creation of DeConvL. The UNetL-based NeuralFuse draws inspiration from Ronneberger et al. (2015), known for its robust performance in image segmentation, and thus, we incorporated it as one of our architectures. Lastly, ConvS, DeConvS, and UNetS are scaled-down versions of the model designed to reduce computational costs and total parameters. The architectures of Convolutional-based and Deconvolutional-based are shown in Table 4, and the architecture of UNet-based generators is in Table 5. For the abbreviation used in the table, ConvBlock means the Convolution block, Conv means a single Convolution layer, DeConvBlock means the Deconvolution block, DeConv means a single Deconvolution layer, and BN means a Batch Normalization layer. We use , , and Adam optimizer. For CIFAR-10, GTSRB, and CIFAR-100, we set batch size for each base model. For ImageNet-10, we set for ResNet18, ResNet50 and VGG11, and for both VGG16 and VGG19.
@@
Appendix C NeuralFuse’s Energy/Accuracy Tradeoff
@@
In Table 6, we report the total weight memory access (T.W.M.A.) using SCALE-SIM. We then showed the energy/accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 5.
@@
We can then use the MAC values to further approximate the end-to-end energy consumption of the whole model. Assume that all values are stored on SRAM and that a MAC represents single memory access. The corresponding MACs-based energy saving percentage (MAC-ES, %) can be derived from Eq. 7 (c.f. Sec. 4.4), and results can be found in Table 8. We can observe that most combinations can save a large amount of energy, except that VGG11 with two larger NeuralFuse (ConvL and DeConvL) may increase the total energy. These results are consistent with the results reported in Table 2. In addition, we also showed the MACs-based energy/accuracy tradeoff between all of the combinations of NeuralFuse and base models under a 1% of bit error rate in Figure 6.
ConvL ConvS DeConvL DeConvS UNetL UNetS ResNet18 16.2 28.7 19.0 26.6 23.2 28.7 ResNet50 24.5 29.8 25.7 28.9 27.4 29.8 VGG11 -21.8 23.9 -11.5 16.0 3.6 23.7 VGG16 5.0 27.3 10.0 23.5 17.4 27.2 VGG19 10.4 28.0 14.4 25.0 20.2 28.0
@@ Appendix D Inference Latency of NeuralFuse
In Table 9, we report the latency (batch_size=1, CIFAR-10/ImageNet-10 testing dataset) of utilizing the different NeuralFuse generators with two different base models, ResNet18 and VGG19. While NeuralFuse contributes some additional latency, we consider this an unavoidable tradeoff necessary to achieve reduced energy consumption within our framework. Although the primary focus of this paper is not on latency, we acknowledge its importance. Future research could explore designing a more lightweight version of the NeuralFuse module or applying model compression techniques to minimize latency. Additionally, we recognize that running NeuralFuse on a general-purpose CPU could lead to different latency and energy consumption figures due to various influencing factors like CPU architecture and manufacturing processes.
@@ Appendix E Ablation Studies
Here, we study the effect of used in EOPM (Eq. 5). In Figure 7, we report the results for ConvL and ConvS on CIFAR-10 pre-trained ResNet18, under a 1% bit error rate (BER). The results demonstrate that if we apply larger , the performance increases until convergence. Specifically, for ConvL (Figure 7(a)), larger empirically has a smaller standard deviation; this means larger gives better stability but at the cost of time-consuming training. In contrast, for the small generator ConvS (Figure 7(b)), we can find that the standard deviation is still large even trained by larger ; the reason might be that small generators are not as capable of learning representations as larger ones. Therefore, there exists a tradeoff between the stability of the generator performance and the total training time. In our implementation, choosing or is a good balance.
Tradeoff Between Clean Accuracy (CA) and Perturbed Accuracy (PA).
@@
| Base Model | CA | PA | ConvL | |||
| CA (NF) | PA (NF) | RP | ||||
@@
Note. CA (%): clean accuracy; PA (%): perturbed accuracy; NF: NeuralFuse; and RP: total recover percentage of PA (NF) vs. PA
@@
Note. BER: the bit-error rate of the base model; CA (): clean accuracy; PA (): perturbed accuracy; UIP: universal input transformation parameter; and RP: total recovery percentage of PA (UIP) vs. PA
@@ Appendix F Additional Experimental Results on Relaxed Access Settings
We conducted more experiments on Relaxed Access settings to show that our NeuralFuse can protect the models under different BER The results can be found in Sec. F.1 (CIFAR-10), Sec. F.2 (GTSRB), Sec. F.3 (ImageNet-10), and Sec. F.4 (CIFAR-100). For comparison, we also visualize the experimental results in the figures below each table.
@@ F.1 CIFAR-10
Base Model NF CA 1% BER 0.5% BER PA CA (NF) PA (NF) RP PA CA (NF) PA (NF) RP ResNet18 ConvL 92.6 89.8 87.8 1.7 48.8 90.4 87.9 2.2 17.8 ConvS 88.2 59.5 11 20.6 91.7 78.4 8.3 8.3 DeConvL 38.9 89.6 88.5 0.8 49.6 70.1 90.2 90.0 0.2 19.9 DeConvS 12.4 82.9 68.8 6.4 29.9 11.6 84.1 79.9 3.6 9.8 UNetL 86.6 84.6 0.8 45.6 89.7 86.3 2.4 16.2 UNetS 84.4 68.8 6.0 29.8 90.9 80.7 5.8 10.7 ResNet50 ConvL 92.6 85.5 53.2 22 27.1 90.3 86.5 3.2 25.5 ConvS 85.2 34.6 14 8.5 90.8 73.3 8.7 12.3 DeConvL 26.1 87.4 63.3 21 37.2 61.0 89.5 87.2 2.5 26.2 DeConvS 9.4 82.4 42.2 17 16.1 10.3 90.3 75.5 8.1 14.5 UNetL 86.2 75.5 12 49.4 89.9 83.9 3.6 22.9 UNetS 77.3 56.2 19 30.1 89.7 76.1 7.2 15.1 VGG11 ConvL 88.4 89.6 87.2 2.9 45.1 89.8 87.0 1.3 23.3 ConvS 84.9 66.3 7.5 24.1 88.2 74.5 5.7 10.9 DeConvL 42.2 89.3 87.2 2.6 45.0 63.6 89.6 86.9 1.1 23.2 DeConvS 11.6 85.6 68.2 7.1 26.0 9.3 88.3 75.7 4.6 12.1 UNetL 87.1 83.6 1.3 41.4 88.0 82.4 1.8 18.8 UNetS 85.5 72.7 4.6 30.5 88.1 75.8 4.3 12.2 VGG16 ConvL 90.3 90.1 86.0 6.2 50.3 90.2 88.5 0.9 21.9 ConvS 87.4 59.6 12 23.9 89.9 77.8 4.8 11.1 DeConvL 35.7 89.7 85.5 6.8 49.8 66.6 89.7 88.2 1.0 21.4 DeConvS 7.9 86.8 66.5 11 30.8 8.1 90.0 78.4 4.7 11.8 UNetL 87.4 83.4 4.4 47.7 89.0 86.2 1.5 19.6 UNetS 87.4 71.2 8.2 35.5 89.0 80.2 3.5 13.7 VGG19 ConvL 90.5 89.8 77.7 19 41.7 90.4 88.1 1.8 23.9 ConvS 87.3 52.7 17 16.7 89.6 74.5 9.0 10.3 DeConvL 36.0 86.3 78.4 18 42.4 64.2 90.4 88.5 1.4 24.3 DeConvS 12.0 86.5 58.2 18 22.2 12.4 89.7 75.2 8.6 11.0 UNetL 86.3 82.1 4.8 46.0 89.1 85.0 2.7 20.8 UNetS 86.3 66.4 13 30.4 89.2 77.1 7.3 12.9 • Note. CA (): clean accuracy; PA (): perturbed accuracy; NF: NeuralFuse; and RP: total recovery percentage of PA (NF) vs. PA